# Canvil: Designerly Adaptation for LLM-Powered User Experiences

### v1

@@ 4.3. User Interface
The Canvil interface resembles an interactive card. The card itself is separated into two areas: the Main Form and Playground Area. When a user selects a Canvil, a property menu is invoked that can open up additional panels for styling, response generation, and settings. Canvils can be freely placed on and moved around the Figma canvas, allowing model tinkering to take place in close proximity to relevant designs. Canvils can also interact directly with designs by reading inputs from and writing model outputs to their text layers. We detail each of Canvil’s features in this section and connect them with our design goals.
Figure 1. An overview of the Canvil interface. A: A blank Canvil with its property menu invoked. The property menu has options for styling (1), model response generation (2), and model settings configuration (3). The Main Form (4) contains text input fields for structured authoring of system prompts. The Playground Area (5) allows users to quickly test inputs and view model outputs. B: A Canvil with some copies styled with pre-set color options in the property menu. Like any other native Figma object, Canvil allows for stateful duplication. C: Canvil supports collaboration by default. Here, we see Figma users collaboratively authoring a Canvil titled “Yoda impersonator.”
On the very left, a white card that resembles a form (the Main Form) with a section at the bottom in light gray (the Playground Area). A dark gray property menu is on top of the card, with 3 options: a color picker, Generate, and Settings. In the middle, the same card is in front while cards of different colors are fanned out behind it, while 6 color options are open in the property menu. On the very right, a green card titled Yoda impersonator with text populated in each of its fields is collaborative edited by multiple cursors.
Figure 1. An overview of the Canvil interface. A: A blank Canvil with its property menu invoked. The property menu has options for styling (1), model response generation (2), and model settings configuration (3). The Main Form (4) contains text input fields for structured authoring of system prompts. The Playground Area (5) allows users to quickly test inputs and view model outputs. B: A Canvil with some copies styled with pre-set color options in the property menu. Like any other native Figma object, Canvil allows for stateful duplication. C: Canvil supports collaboration by default. Here, we see Figma users collaboratively authoring a Canvil titled “Yoda impersonator.”
@@ 4.3.1. Main Form
The Main Form provides a means of authoring a system prompt (DG1) in a structured manner via a multi-field form. Unlike prior systems that offer more or less an open text field for system prompting (OpenAI, 2023a; Petridis et al., 2023b, a), we chose to enforce structure because it provides mental scaffolding to reason about the system prompt in a UX context from multiple facets (e.g., with whom will the LLM interact, and how should the LLM meet their goals?), thus providing more opportunities for designers to integrate user context into the LLM (DG3). Additionally, designers can collaboratively author system prompts (DG4) by each working on a particular field and collectively deliberating afterwards. Breaking down the system prompt into smaller units also allows for more fine-grained experimentation (DG5)—designers can copy a Canvil and tweak a specific field to compare how that change impacts model behavior.
Below are the fields, with a brief description of each, in the Main Form. The field titles (in bold) are always visible on the interface, while the field descriptions can be accessed by hovering over an info icon beside each field title.
@@
We derived these fields from synthesizing system prompting guidelines offered by technical tutorials (OpenAI, 2023c; Microsoft, 2023) as well as NLP literature (Deshpande et al., 2023). We note that our fields are just one possible way to structure a system prompt, and there may be some overlaps between the fields. As discussed in Section 2.2, no “gold standard” currently exists for prompt authoring, so we chose to follow the current recommended guidelines when designing the Main Form.
@@ 4.3.2. Playground Area
Below the Main Form, we provide a Playground Area as an easy way to send user inputs to the model and test model responses (DG5) within Canvil. The designer may test on the same Canvil multiple times or duplicate one with its entire state and test different tweaked versions (Fig. 1B). Designers are likely to find stateful duplication intuitive as it is available on all native objects in the Figma canvas (DG2).
@@ 4.3.3. The Generate Panel
To generate a response from a model adapted with the system prompt authored on the Main Form, the user selects the “Generate” option from Canvil’s property menu, which takes them to the Generate Panel. The panel has two modes: Playground and Design. Both modes contain an option to copy the raw prompt to one’s clipboard so it can be tested in a separate environment (e.g., a Python notebook during collaboration with a data scientist), if desired.
@@
Design Mode. The Design Mode (Fig. 2B) is invoked when selecting “Using design” from the dropdown. In this mode, Canvil navigates the design layers on the user’s Figma canvas, reading text inputs from a specified layer(s) from those designs, and writing model responses to a specified layer(s). This mode leverages the hierarchical layer structure of Figma designs (DG2) and implements the “input-output” LLM-interaction proposed by Petridis et al. (Petridis et al., 2023a).
The user can directly select a layer from their Figma canvas and use the “Set read layer(s)” button to bind it as Canvil’s input. This allows Canvil to retrieve and use the text from the bound layer as user input. If the user has inputs from matching layers [^6] We consider two layers as “matching” if they both have the same name and are of the same Figma layer type. Matching layers are common in designs with related information split up across different UI components. For example, a chat UI may have many matching layers that collectively contain one user’s dialogue across multiple conversation rounds. from the same screen, they can check the option to include them as input, and Canvil will read from those layers accordingly. Similarly, the user can select a layer to which the model writes its response. However, when implementing the checkbox for writing to multiple layers, we realized that writing the same output to multiple locations on the same screen is rarely meaningful for designers. As such, we designed the write layer checkbox to search for matching layers across screens for simultaneous updates in multiple screens along a user flow, ensuring content consistency within a particular matching layer.
@@ 4.3.4. The Settings Panel
The Settings panel (Fig. 2C), accessible through the “Settings” option from Canvil’s property menu, contains some basic model settings for the user to configure, along with a field for the user to enter their OpenAI API key required for response generation. We distilled the list of settings in the OpenAI API to four (model selection, temperature [^7] We rename “temperature” to “creative randomness” to provide a more descriptive name for those not as familiar with LLMs (confusion over the temperature parameter was raised in our formative study—see Section 3.2.3)., maximum generation length, stop words) that may be useful to non-AI experts such as designers. The “Update Settings” button saves the settings to Canvil’s state, which is preserved when the Canvil is duplicated.
@@
From left to right: a Canvil with a property menu above it; a panel with an option to generate a model response after reading the playground input; a panel with an option to generate a model response after setting design layers to read the input from and write the output to; a panel with a model selector and model settings.
@@ 4.4. Implementation
Canvil is a Figma widget [^8] https://www.figma.com/widget-docs/ implemented in TypeScript using the Figma API. [^9] https://www.figma.com/developers/api A Figma widget differs from a Figma plugin [^10] https://www.figma.com/plugin-docs/ by its collaborative nature. A widget is available to all users of the canvas that it is placed in, and maintains a common state that supports multi-user editing by default. Users can interact with widgets just like any other native object on the Figma canvas, including moving, duplicating, and styling. In contrast, Plugins are local to an individual user. The collaborative affordances of widgets align well with DG4, so we implemented Canvil as a widget.
Upon the user selecting ”Generate,” Canvil prepares a system prompt using the text written on the main form. Each field in the form is converted to markdown format and is packaged up in ChatLM [^11] https://learn.microsoft.com/en-us/azure/ai-services/openai/how-to/chatgpt for added prompt parsability. Our prompt template is available in Appendix B. The populated template is sent as a system prompt to an OpenAI Chat API endpoint with settings specified in Canvil’s settings panel. If generation was triggered in Playground Mode, any text in the Playground Area will be sent as user input to the model. If generation was triggered in Design Mode, Canvil searches for the read layer(s) specified in Design mode on the user’s current Figma canvas, retrieves the text within those layers, and sends them off to the API endpoint as user input.
@@ 5. Design Study
With Canvil as a technology probe, we conducted a task-based design study with 17 participants organized into 6 groups. The study was conducted on a per-group basis; each session was 90 minutes in length. Our study investigated the following research questions:
@@
    Implications of adaptation: How does engagement with designerly adaptation impact design workflows and outputs in the context of our study?
@@
    Adaptation at large: What are designers’ attitudes toward designerly adaptation beyond the context of our study?
@@ 5.1. Participants
We recruited our participants from two channels. First, we distributed study invites to designers via email, professional interest groups, and word of mouth at a large technology company. Some of these invitees took part in our formative study. We also snowball sampled by asking our invitees to refer us to other designers they work with who may also be interested in participating. From this channel, we only selected designers with prior experience working on LLM-powered products and features. Second, we sent study invites to a Slack workspace for HCI and design maintained by a large, public academic institution in the United States. The population in the Slack consists primarily of design students and early-career designers. From this channel, we only selected those without any experience working on LLM-powered products and features. Our goal of recruiting from these two channels was to capture potential disparities that may arise in our findings hinging on prior experience working with LLMs.
Invitees first filled out a screening form with information on demographics, professional background, prior exposure to LLMs, and group member preferences. We then organized all candidates from the technology company into groups based on member preferences and availabilities, and did the same for candidates from the academic institution. We left recruitment open for both channels as we conducted the studies and stopped when we reached data saturation. In the end, we had 6 groups (3 groups from each channel) with 17 participants (8 from the technology company and 9 from the academic institution). Five groups were of size 3 and one of size 2. All participants were based in the United States. Participant and group details can be found in Table 2 in Appendix C.
All study sessions were recorded and transcribed. We sent each participant a \$75 USD gift card as an honorarium after the study. This study was approved by the Institutional Review Boards of all organizations involved.
@@ 5.2.1. Setup
We assembled these instructions cognizant of the fact that this was a time-constrained study, and we wanted participants to specifically focus on the area of the interface where users interact with an LLM. To reduce design overhead unrelated to our research questions, we provided starter UIs for participants to work with, along with basic UI components such as text, buttons, and sticky notes (Fig. 3B). If participants had an idea but did not have time to execute it, we encouraged them to describe it on a sticky note next to their designs.
To better probe research-informed model adaptation—that is, adaptation across varying user contexts informed by user research (DG3)—we created descriptions for three hypothetical user groups of this new Feasto feature based on users’ geographic region: west coast of the U.S., Turkey, and India. These descriptions are not meant to act as user personas, but rather high-level sketches of the customs and preferences that may be prevalent in the region. We crafted the user group descriptions to vary along three key dimensions (assuming everyone used Feasto in English): dietary restrictions, access to ingredients, and menu style preferences. Definitions for these dimensions, along with the full user group descriptions, are available in Appendix D.
@@ 6. Results
Our quantitative analysis showed that Canvil had a mean SUS score of 69.94 (std = 12.18), meaning its usability was “above average” (usability.gov, 2022). In this section, however, we focus on the qualitative insights on designerly adaptation from our design study. We specifically showcase the implications of satisfying Canvil’s design goals (Section 4.1) on design workflows and outputs. Our results from designers without prior experience working with LLMs did not differ noticeably from those with experience, except that the latter drew more connections between adaptation and their past model tinkering workflows. We present text that participants wrote in Canvil as purple and italicized.
6.1. Authoring System Prompts as Part of Design Workflows
6.1.1. Research-Informed System Prompting as a Design Activity (DG1, DG3)
@@
When authoring system prompts in Canvil (DG1), designers frequently referenced the provided user research documents. It is through this process that they embedded user context—e.g., varying lifestyles, customs, preferences—into system prompts (DG3). For example, in P2’s Canvil for the American user group, they wrote under Audience Setting that the model is serving users who like diverse cuisine but care about the environment and sustainability and are health-conscious. Designers also adjusted their prompts between user groups in attempts to create more customized and meaningful user experiences. In the Canvil for the user group from India, P4 wrote that The model should use Indian terminologies to describe the recipe. Eg: eggplant is brinjal in India. P5 experimented with model variations within one user group, choosing distinctly different personalities for each model. For their Indian user group, they populated one Canvil’s Model Profile field with The model is a head chef of a 5 star restaurant situated in New Delhi. It is a busy day for the restaurant and the chef is low on time. and another with The model is a mother who is helping her son cook quick meals in hostel. In the end, they noted that “the models did very well” in taking on these different profiles.
@@
In particular, designers used the Guardrails field in Canvil to address potential violations of users’ dietary preferences and restrictions, along with model misuse. For example, P12 used information provided on the Turkish user group to create the following guardrail: The model should not include recipes with pork or alcohol and must respect the fasting period of Ramadan by suggesting suitable pre-dawn and post-dusk meals. A few participants noted that some user groups may require more strict model guardrails than others. P15 identified a subtle but key requirement between two user groups while authoring their Canvils:
@@
Designers sought to address user needs through combined efforts in prompt authoring and UI design. For example, P15 provided alternatives via buttons to support more flexible user input of ingredients going into the model. P8 described their UI design in response to observing the limits of what can be achieved via writing system prompts: “Sometimes even in the instructions that I gave to Canvil, it wasn’t really reflecting that [desired] granularity until I pushed it further. In my design I ended up putting a little textbox area where people can specify how detailed they want the instructions.” Fig. 4 contains some examples of designers’ UIs from our study.
In short, the structure provided by prompt authoring interface of Canvil is compatible with designers’ reasoning about user needs and preferences based on user research. Canvil supports this reasoning process to inform both the system prompts and UI design in an integrated fashion, and makes authoring system prompts a natural part of design activities.
Figure 4. Examples of UIs designers created for Feasto’s 3-course meal planner during the study.
Figure 4. Examples of UIs designers created for Feasto’s 3-course meal planner during the study.
6.1.2. Tinkering Allowed for Deeper Reasoning of User-LLM Interaction (DG2, DG5)
Canvil’s integration in the design environment (DG2) allowed designers to consider model behavior and user interaction simultaneously. Many designers iterated multiple times on their system prompts (DG5) until they were satisfied about the adapted model. P3 reformatted the output to better suit their existing UI: “the first run was a paragraph that I felt like was really hard to read, so I added some core instructions saying give me a numbered list.” P13 shared that their model was repeating the same ingredients for each course, so they “tweaked the level of detail and the recipes to add more variety.” P11 discovered that, in general, “stating [instructions] in the positive instead of the negative” yielded better results for their envisioned experience.
Designers’ tinkering with the adapted model not only helped them iterate on their system prompts, but also allowed them to more deeply reason about user interactions with LLMs and reflect on the whys of UI design. For P14, grappling with the Example Inputs and Outputs field in Canvil prompted reasoning about the optimal interaction pattern to design with:
“The example inputs and outputs were a little bit hard for me to think about, cause I was wondering for the output, is it just the recipe? Or would it be more of a conversation, like oh, do you have dietary restrictions? Do you have any XYZ preferences? Like there’s some logic that goes into asking those questions and I was wondering which one should I think about first versus later.” [P14]
Tinkering with the adapted model often led to evolving of or experimenting with UI ideas. P17 considered different interaction patterns for their interface—a chat-based interface versus a form-filling GUI—after discovering how sensitive LLMs can be to prompts. P4 mentioned that they started off with “a specific user base in mind and I wanted a specific tone,” but the design of the UI became clearer after they saw an ideal output after multiple iterations: “My design direction changed when I saw the output. I iterated multiple times and what I like the best about this [response] was that it was short and [ideal] for scrolling in a screen. I tried that out [in my new UI] and it worked pretty well.” Examples of quoted participants’ designs can be found in Fig. 4.
6.2. Designers Envisioned Collaborative Workflows with Designerly Adaptation
6.2.1. Collaborative Uses of Canvil in Design Teams (DG4)
The design task in our study was an individual activity because we expected collaborative efforts to begin after designers had made some individual attempts at adaptation and designed with the adapted model (Feng et al., 2023c). Indeed, when reflecting on Canvil’s collaborative capabilities (DG4), designers thought the workflow matched their own experiences. P8 did not consider a “divide and conquer” authoring approach—where each team member is responsible for one Canvil field—as necessary unless the team was under a tight time constraint, and instead saw more potential in a “tinker and unite” approach: “Since it doesn’t take too much time to generate a response, I feel like I would lean towards, let’s all try different things and see how we can use this tool to get the best response.” P8 added that in order for this approach to work, some guidelines should be set for collaborators “so that all of our responses are consistent, like everyone has the same Audience Setting and [uses] the same kind of wording.” P4 agreed and suggested initial brainstorming sessions for collaborators so that they could “find a way to come to common grounds for all these [fields].” P6, after hearing P4’s suggestion, noted that using Canvil within FigJam may be useful “for ideation or brainstorming” before bringing more polished Canvils into Figma.
On the other hand, some saw potential in the “divide and conquer” approach, especially when expertise in a design team may be distributed. P7, who specialized more in visual design, said that they were “struggling a little bit with like how to prioritize all the content, and I feel like that would be specialties a content designer would bring in.” P8, who initially preferred to “tinker and unite,” agreed that it would be valuable to also leverage distributed expertise: “In a design team, there might be someone who knows a user persona the best, and they can be assigned to Audience Setting. Then you have content designers on maybe example inputs and outputs.”
6.2.2. Collaboration and Knowledge Sharing Afforded by Canvil (DG2, DG4)
Many designers commented on how Canvil, as a widget integrated into Figma’s design environment (DG2), afforded collaboration (DG4). P5 appreciated that they could leverage Figma’s built-in collaboration features: “If I see my fellow designer’s Canvil, and if I want to change something then and there, I can drop a comment on it just like a normal component in Figma, which we’ve been doing in our everyday design work.” P10 shared that their team relies on a spreadsheet to keep track of system prompt iterations and working in a canvas environment would be a significant improvement.
In particular, participants saw potential for more effective knowledge sharing about adaptation with Canvil. P10 commented that: “You could have a master Canvil and then you could make copies, and [others] can then do their own interpretations.” P11 agreed and added that seeing iterations on a canvas can help find inspiration in others’ work: “I think anytime you line up different iterations together, you notice: ohh that person, did you know they had that approach? That’s a good idea. And I’m gonna try that over here. I think it really is an aid to experimentation.” We observed other instances of knowledge sharing as well when designers viewed others’ Canvils. For example, P4 noted that they were inclined to consider more dimensions for their system prompt after seeing P6’s work: “[looking at P6’s Canvil] made me thinking about timers and Hindi slang.” Some also realized new capabilities of LLMs by looking at others’ prompts and outputs—P1 shared that “I didn’t realize at first that you can actually make the [LLM] generate multiple recipes, just like what P2 did.” P1’s groupmates (P2 and P3) were also intrigued when they saw that P1 had used one Canvil to generate example inputs for another Canvil to use. These social and collaborative affordances differentiate Canvil from prior systems for empowering designers to tinker with AI (Subramonyam et al., 2021a; Petridis et al., 2023b; Carney et al., 2020).
6.2.3. Collaborating With Non-Design Stakeholders via Canvil (DG1, DG4)
With system prompt authoring as the primary interaction, Canvil (DG1) ensures easy authorship and sharing within a team (DG4), even across domain boundaries. Many designers thought that Canvil was simple and intuitive enough to be used by not just designers. Consequently, P2 saw Canvil as a useful boundary object (Star and Griesemer, 1989): “I can see from a product manager’s perspective that they would love to play around with the prompts, and it would probably help the designer explain their design choices. I feel they would find more common ground because they both used a similar tool.” P6 agreed that product managers (PMs) have perspectives and expertise to contribute to prompt authoring. Designers also saw potential for deeper collaboration with data scientists and other technical stakeholders. P11 commented that the “Copy raw prompt to clipboard” button in Generate Panel (Section 4.3.3) can allow for direct handoff of system prompts: “[The button] is a way of exporting that so a data scientist can come along and say sure, let me plug that into code.” Even if designers were unable to achieve the desired model behavior themselves, P17 thought Canvil was helpful in specifying desired changes: “If I want the model to respond in this way versus that way, just having something tangible to show engineering partners where the tweak would need to be, would be helpful. And I think [that’s] obviously easier with Canvil.” P15 believed that rise of natural language as a shared representation can allow for more transparent and balanced information flow between disciplinary boundaries:
“I think on the designer side, there is apprehension about [AI] being in different skill set. The [technical] teams just don’t really embrace designers, cause this is ‘their’ skill set. They’re gonna do it quicker and better, and that’s probably true in almost all circumstances except natural language. And except when it comes to an intuitive experience. So it’s about making the teams meet each other as opposed to one team being more reticent than the other.” [P15]
Designers also acknowledged that designerly adaptation may only be one piece of the larger puzzle to steer an LLM’s behavior and integrate it into a product. Some desired changes, such as factual grounding and connecting to external knowledge bases, may still need to be left to technical teams. P12 even saw factual grounding as a prerequisite to designerly adaptation: “because everything hinges on you being able to get the model specificity correct.” Both P12 and P14 wanted Canvil to be able to connect to other models built and maintained by their own teams, as the behavior of unadapted OpenAI models may be too generic to use as an effective design material. P14 comments: “I think linking to the vanilla ChatGPT model isn’t specific enough. If there is a way to somehow have Canvil link to the devs’ [models] while I’m designing and I’m able to see what type of experience and what type of response our users are getting, I think that would be super helpful.”
While we currently offer only three OpenAI models to use with Canvil (GPT-3.5-Turbo, GPT-3.5-Turbo-16k, GPT-4), more may be added with lightweight engineering as the widget is communicating with the model via API endpoints. As such, we can envision future support for more model families, including teams’ custom models.
6.3. Designerly Adaptation at Large
6.3.1. Integrating Adaptation into Design Practice

### v2

@@ 4.2. User Interface
The Canvil interface resembles an interactive card. The card itself is separated into two areas: the Main Form and Playground Area. When a user selects a Canvil, a property menu is invoked that can open up additional panels for styling, response generation, and settings. Canvils can be freely placed on and moved around the Figma canvas (DG1), allowing model tinkering to take place in close proximity to relevant designs. Canvils can also interact directly with designs by reading inputs from and writing model outputs to their text layers. We detail each of Canvil’s features in this section and connect them with our design goals.
Figure 2. An overview of the Canvil interface. A: A blank Canvil with its property menu invoked, with some additional copies styled with pre-set color options. The property menu has options for styling (1), model response generation (2), and model settings configuration (3). The Main Form (4) contains text input fields for structured authoring of model behavioral specifications. The Playground Area (5) allows users to quickly test inputs and view model outputs. B: Canvil supports collaboration by default. Here, we see Figma users collaboratively authoring a Canvil titled “Yoda impersonator.” Like any other native Figma object, Canvil also supports stateful duplication.
[…]
On the left, a white card that resembles a form (the Main Form) with a section at the bottom in light gray (the Playground Area). A dark gray property menu is on top of the card, with 3 options: a color picker, Generate, and Settings. On the right, a green card titled Yoda impersonator with text populated in each of its fields is collaborative edited by multiple cursors.
@@ 4.2.1. Main Form
The Main Form provides a structured means of authoring model behavioral specifications in natural language via a multi-field form. Unlike prior systems that offer an open text field for defining similar specifications (OpenAI, 2023a; Petridis et al., 2023b, a), we chose to enforce structure because it provides mental scaffolding to reason about model behavior in a UX context from multiple facets (e.g., with whom will the LLM interact, and how should the LLM meet their goals?), thus providing more opportunities for designers to translate design requirements into behavioral constraints for the LLM (DG2). Breaking down specifications into smaller units also allows for more fine-grained tinkering and iteration (DG3)—designers can copy a Canvil and tweak a specific field to compare how that change impacts model behavior. Additionally, because Figma widgets support multi-user interaction by default, designers can collaboratively author specifications (DG4).
Below are the fields, with a brief description of each, in the Main Form. The field titles, in bold, are always visible on the interface, while the field descriptions can be accessed by hovering over an info icon beside each field title.
@@
We derived these fields from synthesizing best practices from technical tutorials and documentation (OpenAI, 2023c; Microsoft, 2023; Anthropic, 2024) as well as NLP literature (Deshpande et al., 2023). We note that our fields are just one possible way to translate design requirements into model behavioral specifications, and there may be some overlaps between the fields. As discussed in Section 2.2, no “gold standard” currently exists for this translation, so we chose to follow the current recommended best practices when designing the Main Form.
@@ 4.2.2. Playground Area
Below the Main Form, we provide a Playground Area as an easy way to send user inputs to the model and test model responses (DG3) within Canvil. The designer may test on the same Canvil multiple times or duplicate one with its entire state and test different tweaked versions (Fig. 2B). Designers are likely to find stateful duplication intuitive as it is available on all native objects in the Figma canvas (DG1).
@@ 4.2.3. The Generate Panel
To generate a response from a model adapted with specifications from the Main Form, the user selects the “Generate” option from Canvil’s property menu, which takes them to the Generate Panel. The panel has two modes: Playground and Design. Both modes contain an option to copy the raw prompt to one’s clipboard so it can be tested in a separate environment (e.g., a Python notebook during collaboration with a data scientist), if desired.
@@
Design Mode. The Design Mode (Fig. 3B) is invoked when selecting “Using design” from the dropdown. In this mode, Canvil navigates the design layers on the user’s Figma canvas, reading text inputs from a specified layer(s) from those designs, and writing model responses to a specified layer(s). This mode leverages the hierarchical layer structure of Figma designs (DG1) and implements the “input-output” LLM-interaction proposed by Petridis et al. (Petridis et al., 2023a).
@@ 4.2.4. The Settings Panel
The Settings panel (Fig. 3C), accessible through the “Settings” option from Canvil’s property menu, contains some basic model configuration settings. We distilled the list of settings in the OpenAI API to four (model selection, temperature, [^5] We rename “temperature” to “creative randomness” to provide a more descriptive name for those not as familiar with LLMs (confusion over the temperature parameter was raised in our formative study—see 3.2.3). maximum generation length, stop words) that may be useful to non-AI experts such as designers. The “Update Settings” button saves the settings to Canvil’s state, which is preserved when the Canvil is duplicated.
@@
[…]
From left to right: a panel with an option to generate a model response after reading the playground input; a panel with an option to generate a model response after setting design layers to read the input from and write the output to; a panel with a model selector and model settings.
@@ 4.3. Implementation
Canvil is a Figma widget [^6] https://www.figma.com/widget-docs/ implemented in TypeScript using the Figma API [^7] https://www.figma.com/developers/api. A Figma widget differs from a Figma plugin [^8] https://www.figma.com/plugin-docs/ by its collaborative nature. A widget is available to all users of the canvas that it is placed in, and maintains a common state that supports multi-user editing out-of-the-box. Users can interact with widgets just like any other native object on the Figma canvas, including moving, duplicating, and styling. In contrast, Plugins are local to an individual user. The collaborative affordances of widgets align well with DG4, so we implemented Canvil as a widget.
Upon the user selecting ”Generate,” Canvil prepares a system prompt using specifications from the Main Form. Each field in the form is converted to markdown format and is packaged up in ChatLM [^9] https://learn.microsoft.com/en-us/azure/ai-services/openai/how-to/chatgpt for added prompt parsability. Our prompt template is available in Appendix B. The populated template is sent to an OpenAI Chat API endpoint with settings specified in Canvil’s settings panel. If generation was triggered in Playground Mode, any text in the Playground Area will be sent as user input to the model. If generation was triggered in Design Mode, Canvil searches for the read layer(s) specified in Design mode on the user’s current Figma canvas, retrieves the text within those layers, and sends them off to the API endpoint as user input.
@@ 5. Design Study
With Canvil as a technology probe, we conducted a task-based design study with 17 participants organized into 6 groups. The study was conducted on a per-group basis; each session was 90 minutes in length. Our study investigated the following two research questions:
@@
    A new design code in practice: How do designers make use of the two-way translation between design requirements and UX designs enabled by designerly adaptation via Canvil, within the context of our study?
@@
    Designerly adaptation at large: What are designers’ attitudes toward designerly adaptation beyond the context of our study?
@@ 5.1. Participants
We invited participants to our study via email, professional interest groups, Slack channels for designers run by a public U.S. university, and word of mouth at a technology company. Full recruitment details can be found in Appendix C. Invitees first filled out a screening form with information on demographics, professional background, prior exposure to LLMs, and group member preferences. We selected our participants on a first-come first-serve basis, but accounted for their prior experiences designing LLM-powered UX—we aimed for only half our participant pool to have such experience. We then organized all candidates into groups based on member preferences and availabilities. We left recruitment open as we conducted the studies and stopped when we reached data saturation. In the end, we had 6 groups with 17 participants total; five groups were of size 3 and one of size 2. All participants were based in the United States. Participant and group details can be found in Table 2 in Appendix D.
All study sessions were recorded and transcribed. We gifted each participant a \$75 USD gift card after the study. This study was approved by the IRBs of all organizations involved.
@@ 5.2.1. Setup
As part of the exercise, we intentionally left designers to determine how exactly the LLM powers this feature. We also assembled these instructions cognizant of the fact that this was a time-constrained study, and we wanted participants to specifically focus on the area of the interface where users interact with an LLM. To reduce design overhead unrelated to our research questions, we provided starter UIs for participants to work with, along with basic UI components such as text, buttons, and sticky notes (Fig. 4B). If participants had an idea but did not have time to execute it, we encouraged them to describe it on a sticky note next to their designs. Finally, we picked a universally understood domain (food and meal planning) to avoid any variance in domain-specific knowledge.
To better probe research-informed model adaptation—that is, adaptation across varying user contexts informed by user research (DG3)— we created descriptions for three hypothetical user groups of this new Feasto feature based on users’ geographic region: west coast of the U.S., Turkey, and India. These descriptions are not meant to act as user personas, but rather high-level sketches of the customs and preferences that may be prevalent in the region. We crafted the user group descriptions to vary along three key dimensions (assuming everyone used Feasto in English): dietary restrictions, access to ingredients, and menu style preferences. Definitions for these dimensions, along with the full user group descriptions, are available in Appendix E.
@@ 6. Results
Our quantitative analysis showed that Canvil had a mean SUS score of 69.94 (std = 12.18), meaning its usability was “above average” (usability.gov, 2022). In this section, however, we focus on the qualitative insights from our design study that shed light on our two research questions. We present text that participants wrote in Canvil as purple and italicized.
6.1. RQ1: A New Design Code in Practice
6.1.1. From Design Requirements to UX Designs
Designerly adaptation, as a design code, is meant to facilitate translation of design requirements to concrete designed objects—in our case, UX designs. We saw robust evidence of designers effectively leveraging this translation in our study. Canvil’s integration into the Figma design environment (DG1) and its support for quick iteration (DG3) allowed designers to more deeply reason about user-LLM interactions based on model behavior shaped by their design requirements. P17, for example, tried two different interaction patterns for their interface—a chat-based interface versus a form-filling GUI—to see which one better handles LLMs’ high sensitivity to their specifications’ wording. P15 provided alternatives via buttons to support more flexible user input of ingredients to the model, while P8 described their UI design in response to observing the limits of what can be achieved via editing the Main Form: “Sometimes even in the instructions that I gave to Canvil, it wasn’t really reflecting that [desired] granularity until I pushed it further. In my design I ended up putting a little textbox area where people can specify how detailed they want the instructions.” Examples of quoted participants’ designs can be found in Fig. 5.
Figure 5. Examples of UI screens designers created for Feasto’s 3-course meal planner during the study.
[…]
6.1.2. From UX Designs to Design Requirements
Design codes facilitate translation from not just design requirements to designed objects, but also from objects back to requirements (Cross, 1982). Here, we discuss the reverse translation of Section 6.1.1. Some designers were more adamant about adapting the model to fit existing UI. P4 mentioned that they “iterated multiple times” on the Main Form to find the ideal response. They often stopped iterating when they viewed that the LLM output satisfied the UX requirement, such as being concise: “what I like the best about this [response] was that it was short and [ideal] for scrolling in a screen. I tried that out [in my new UI] and it worked pretty well” (see Fig. 5).
@@ 6.1.2. From UX Designs to Design Requirements
Many designers also aimed to maintain a consistent UI across the user groups they were designing for, but adjusted the requirements they typed into Canvil to create meaningful customizations of LLM based on nuanced differences in user needs. This was enabled by Canvil’s main form, which implements DG2. For example, in the Canvil for the user group from India, P4 wrote that The model should use Indian terminologies to describe the recipe. Eg: eggplant is brinjal in India. P5 experimented with different Canvil versions within one user group and assigned distinctly different personalities for each model. For their user group in India, they populated one Canvil’s Model Profile field with The model is a head chef of a 5 star restaurant situated in New Delhi. It is a busy day for the restaurant and the chef is low on time. and another with The model is a mother who is helping her son cook quick meals in hostel. After some experimentation, they noted that “the models did very well” in taking on these different profiles.
@@
Iterating on the Guardrails field in Canvil specifically allowed designers to address potential LLM safety concerns, such as violations of users’ dietary restrictions and model misuse. In addition to stating dietary restrictions of their Turkish user group in the Guardrails field, P12 also specified what the model should do in response: The model should not include recipes with pork or alcohol and must respect the fasting period of Ramadan by suggesting suitable pre-dawn and post-dusk meals. A few designers noted that some user groups may require more strict guardrails than others. P15 identified a subtle but key requirement between the Canvils for the American and Turkish user groups:
@@
UX design was never just about the visual UI—user flows, content strategy, customization heuristics, and many non-UI considerations play vital roles in shaping a user experience. We can see through our findings that this is especially true with LLM-powered UX. Even within the same UI, designers may create entirely different experiences based on their (shifting) design requirements by adapting LLMs via Canvil. Thus, through designerly adaptation, designers can more effectively explore and user LLMs as a design material.
6.2. RQ2: Designerly Adaptation at Large
6.2.1. Designers Were Receptive to Designerly Adaptation but also Recognized its Limits